# Post-detection inference for sequential changepoint localization

### v4

@@ 1 Introduction
$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
@@
For instance, rapid shifts in public sentiment often occur on online platforms: a product failure may trigger a sudden surge of negative reviews, or a major event may abruptly alter the tone of social-media discussions. Detecting when such a shift occurs is vital for understanding emerging issues, tracing public reactions to events, or monitoring product reputation in real time. To illustrate the relevance of our methodology in such settings, we conduct an experiment using the Stanford Sentiment Treebank (SST-2) dataset [31], simulating a stream of text reviews whose sentiment abruptly transitions from predominantly positive to predominantly negative. Accurately localizing the changepoint is important: it can help investigate why the change may have occurred.
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha,$ (1.2)
where $\mathbb{P}_{F_{0},T,F_{1}}$ is the data distribution in (1.1), i.e. having the true changepoint at $T$. But this goal is harder to achieve than may initally appear; we elaborate below.
@@
For algorithms with a bounded ARL, we prove that no set $\mathcal{C}$ can satisfy guarantee (1.2) (for algorithms with bounded PFA, we show how to achieve (1.2) in the Supplementary). This impossibility result (which we prove formally later) intuitively holds because it is meaningless to ask for coverage if the algorithm stopped before the changepoint; indeed, on the event $\{\tau<T\}$, we have no information to rule out the possibility that $T=\infty$, and thus the only way to cover $T$ is to output the set of all natural numbers $\mathbb{N}$. Thus, an arguably more sensible goal is to have
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,$ (1.3)
meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$. This definition has been considered by, for example, [5] and [42, 39], who achieved it (in a certain asymptotic sense) in much more restricted settings than the current investigation.
The aforesaid impossibility result no longer applies to the new goal (1.3). It is important to note that we do not prescribe which $\mathcal{A}$ to use in which situation, but simply enable post-hoc inference on $T$ for any user-chosen $\mathcal{A}$. One can view our techniques as providing a wrapper to convert $\mathcal{A}$ from producing only a stopping time $\tau$ to also producing a confidence set for $T$  [^2] In some sense, our goal can be viewed in an analogous way to conformal inference, which wraps around an arbitrary prediction algorithm to convert a “point prediction” into a prediction set, by running the same algorithm on many hypothetical datasets (for full conformal prediction). However, our methods are not distribution-free, and rely on smartly using simulations to accomplish the goal.. As we survey below, there is a rich literature on designing change detection algorithms $\mathcal{A}$, but a substantial paucity of inference methodology for localizing the change after detection; our focus is on the latter goal, assuming the user has chosen a suitable algorithm $\mathcal{A}$ for detection.
While inference for $T$ is intrinsically interesting and important, a side benefit of confidence sets for $T$ is that one can use these to derive post-hoc confidence sets for unknown parameters of $F_{0}$ and $F_{1}$. To clarify, suppose $\mathcal{P}_{i}=\{F_{\theta}\}_{\theta\in\Theta_{i}};i=0,1$. Our method that delivers (1.3) also delivers confidence sets $\mathcal{C}^{\prime}\subseteq\Theta_{0}\times\Theta_{1}$ such that $\mathbb{P}_{\theta_{0},T,\theta_{1}}((\theta_{0},\theta_{1})\in\mathcal{C}^{\prime}\mid\tau\geq T)\geq 1-\alpha,\forall\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},$ where $\mathbb{P}_{\theta_{0},T,\theta_{1}}$ is the distribution of the sequence $X_{1},\dots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{0}}$, $X_{T},X_{T+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{1}}$. We stress that our confidence sets do not rely on asymptotics and are valid in finite samples.
@@ 1.1 Related work
We are aware of only two directly related works: [5] and [39, Chapter 3]. They also chose to aim for the guarantee (1.3) like us. Both [5, 39] provide asymptotically valid inference after the CUSUM procedure [19] declares a change for known $F_{0}$ and $F_{1}$ in an exponential family; the former only constructs a lower confidence bound for $T$ while the latter also provides a confidence set for $T$. Their approaches have notable methodological and theoretical limitations: (i) they are tailored specifically to the CUSUM procedure, (ii) they assume that the pre- and post-change distributions are both known and they belong to an exponential family, (iii) they only provide asymptotic coverage guarantees (valid when both $T\to\infty$ and ARL $\to\infty$). These constraints limit practical applicability, and their confidence set exhibits undercoverage in our experiments.
In sharp contrast, our proposed confidence sets for $T$ are very general: We make no assumptions on the detection procedure $\mathcal{A}$, and our framework acts as a versatile wrapper that can be layered on top of any sequential change detection algorithm. It is not restricted to exponential families — it works for any known $F_{0}$ and any nonparametric post-change class $\mathcal{P}_{1}$ on any general space (e.g., $\mathbb{R}^{d}$) and offers non-asymptotic coverage guarantees. It also extends to composite pre-change classes under some assumptions. Thus, our framework is the first nonparametric, theoretically sound, and widely applicable practical tool for sequential change localization tasks. We summarize the comparison in Table 1.1.
Table 1.1: Comparison of existing methods and our proposed approach
Pre-, Post- Existing Methods [5, 39] Our Proposed Method
@@
($F_{0},F_{1}$) $\in$ exponential family ($T$, ARL $\to\infty$) (any $\mathcal{A},F_{0},F_{1}$) (Theorem 4.1)
@@
($F_{0},\mathcal{P}_{1}$) (any $\mathcal{A},F_{0},\mathcal{P}_{1}$) (Theorem 3.2)
Unknown, Unknown No existing work Assumption 3.5 Nonasymptotic
($\mathcal{P}_{0},\mathcal{P}_{1}$) on $\mathcal{A},\mathcal{P}_{0}$ (any $\mathcal{P}_{1}$) (Theorem 3.8)
@@
Literature on sequential change detection. In sequential change detection, the goal is to detect that a change has occurred as soon as possible, but not to provide inferential tools for when the change occurred. This methodology has proved crucial in many applications, such as quality control, cybersecurity, medical diagnostics, and so on. Classical methods such as Shewhart control charts [27], the CUSUM procedure [19], and the Shiryaev-Roberts approach [29, 23] assume known pre- and post-change distributions. To address scenarios with unknown post-change distributions, generalized likelihood ratio (GLR) procedures [18, 30] were developed. Comprehensive reviews of these methods are provided by [33]. More recently, kernel-based techniques [9], e-detectors [28], etc., have emerged to handle nonparametric settings.
@@
Offline vs. online change localization. As far as we are aware, most existing approaches to changepoint localization focus on offline settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times. Classical works, e.g., [10] study the asymptotic behavior of the MLE for the changepoint and construct likelihood-ratio-based confidence sets. [3] proposed some nonparametric estimators. [38] developed likelihood ratio-based confidence sets for exponential family distributions, while [6] used bootstrap-based tests for each candidate time to construct nonparametric confidence sets with asymptotic coverage. Recently, [34] studies optimal rates for both changepoint detection and localization. [15] constructs confidence sets by evaluating a local test statistic. To re-emphasize, all these methods are for the offline setting. In the sequential setting, [32, 7, 40, 1] develop point estimators for $T$; but very few works construct confidence sets [5, 39], whose drawbacks were discussed earlier.
@@
Inference on $\theta_{0},\theta_{1}$ after stopping. As with the problem of post-detection inference on the changepoint $T$, there is limited literature on valid estimation or inference for the pre- and post-change parameters $\theta_{0}$ and $\theta_{1}$ post-stopping. [42] constructs asymptotic bias-corrected point estimates and confidence intervals for the post-change mean after a mean change is detected by a CUSUM procedure. [41] generalizes these results for the case when there is a possible change in the variance as well. These methods are also discussed in [39]. However, these approaches share several limitations with the existing post-detection changepoint inference methods, as mentioned earlier: they rely on asymptotics, assume specific parametric forms, and apply only to the CUSUM procedure.
@@ 1.2 Contributions and paper outline
  • We propose a general framework for constructing a confidence set for the changepoint based on inverting tests obtained using a universal threshold, assuming all observations to be independent. We show that it provides $1-\alpha$ conditional coverage for known pre-change and arbitrary (nonparametric) post-change classes under no additional assumption in Theorem 3.2. We also extend it for composite pre-change under a suitable assumption in Theorem 3.8.
  • We propose an alternative method of constructing the tests using adaptive thresholds (which is often tighter than the universal threshold), assuming pre- and post-change distributions to be known, in Algorithm 1. We extend this framework to handle parametric pre- and/or post-change classes in Algorithms 2 and 3. These adaptive methods allow the observations to be dependent. We establish corresponding coverage guarantees in Theorems 4.1, 4.3 and 4.4.
  • As a byproduct of our methods, we derive confidence sets for the unknown pre- and post-change parameters and show their coverage guarantees in Theorem 5.1.
@@
The rest of the paper is organized as follows. In Section 2, we establish some fundamental results showing the impossibility of unconditional coverage and linking conditional confidence sets to hypothesis tests. Section 3 develops a general framework for constructing a confidence set for $T$ based on inverting tests obtained using a universal threshold. In Section 4, we introduce a simulation-based approach for constructing confidence sets using an adaptive threshold in the parametric setting and study their conditional coverage probabilities. In Section 5, we derive confidence sets for the changepoint, as well as the pre- and post-change parameters. Section 6 presents an extensive simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach. This article is concluded in Section 7. All the mathematical proofs, implementation details of some algorithms in the paper, and additional experiments are presented in the Supplementary Material.
@@ 2 Some Fundamental Results
In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$.
Notation. For $t\in\mathbb{N}$, let $\mathbb{P}_{F_{0},t,F_{1}}$ denote joint distribution of the sequence $\{X_{n}\}_{n}$ when $X_{1},\dots,X_{t-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$ and $X_{t},X_{t+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$.  [^3] To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions. However, our framework in Section 4 can actually handle dependent data. Under dependence, $F_{0}$ and $F_{1}$ denotes the joint distributions of $X_{1},\dots,X_{t-1}$ and $X_{t},X_{t+1},\dots$ respectively, and is in fact ambivalent to whether $X_{t}$ is drawn independently of $X_{1},\dots,X_{t-1}$ or conditional on it, so we do not emphasize the distinction in our notation. As a corner cases, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$. In parametric settings, we write $\mathbb{P}_{\theta_{0},t,\theta_{1}}$ and $\mathbb{P}_{\theta_{0},\infty}$ as shorthand for $\mathbb{P}_{F_{\theta_{0}},t,F_{\theta_{1}}}$ and $\mathbb{P}_{F_{\theta_{0}},\infty}$ respectively.
@@
The proposition says that for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$. Thus, for post-detection inference, the usual notion of coverage is not an appropriate one. Instead, we focus on conditional coverage (1.3), a more reasonable criterion, which was also considered in prior work [5, 39]. The key idea behind our methodology is that for each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
@@
$\mathcal{C}:=\{t\in\mathbb{N}:t\leq\tau,H_{0,t}\text{ is not rejected}\}$ (2.2)
@@
Proposition 2.2.
Let $\tau$ be a stopping time and $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying (2.4). Then, the test defined by $\phi^{(t)}(X_{1},\cdots,X_{\tau})=\mathds{1}(t\notin\mathcal{C})$ satisfies
$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.3)
Conversely, given any test $\phi^{(t)}$ for $H_{0,t}$ using data up to $\tau$ satisfying (2.3), the inverted confidence set $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\penalty 10000\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.4)

### v6

@@ 1 Introduction
$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1)
@@
For instance, rapid shifts in public sentiment often occur on online platforms: a product failure may trigger a sudden surge of negative reviews, or a major event may abruptly alter the tone of social-media discussions. Detecting when such a shift occurs is vital for understanding emerging issues, tracing public reactions to events, or monitoring product reputation in real time. To illustrate the relevance of our methodology in such settings, we conduct an experiment using the Stanford Sentiment Treebank (SST-2) dataset [socher2013recursive], simulating a stream of text reviews whose sentiment abruptly transitions from predominantly positive to predominantly negative. Accurately localizing the changepoint is important: it can help investigate why the change may have occurred.
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha,$ (2)
where $\mathbb{P}_{F_{0},T,F_{1}}$ is the data distribution in (1), i.e. having the true changepoint at $T$. But this goal is harder to achieve than may initally appear; we elaborate below.
@@
For algorithms with a bounded ARL, we prove that no set $\mathcal{C}$ can satisfy guarantee (2) (for algorithms with bounded PFA, we show how to achieve (2) in the Supplementary Section F). This impossibility result (which we prove formally later) intuitively holds because it is meaningless to ask for coverage if the algorithm stopped before the changepoint; indeed, on the event $\{\tau<T\}$, we have no information to rule out the possibility that $T=\infty$, and thus the only way to cover $T$ is to output the set of all natural numbers $\mathbb{N}$. Thus, an arguably more sensible goal is to have
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,$ (3)
meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$. This definition has been considered by, for example, [ding2003lower] and [WU20063625, wu2007inference], who achieved it (in a certain asymptotic sense) in much more restricted settings than the current investigation.
The aforesaid impossibility result no longer applies to the new goal (3). It is important to note that we do not prescribe which $\mathcal{A}$ to use in which situation, but simply enable post-hoc inference on $T$ for any user-chosen $\mathcal{A}$. One can view our techniques as providing a wrapper to convert $\mathcal{A}$ from producing only a stopping time $\tau$ to also producing a confidence set for $T$ As we survey below, there is a rich literature on designing change detection algorithms $\mathcal{A}$, but a substantial paucity of inference methodology for localizing the change after detection; our focus is on the latter goal, assuming the user has chosen a suitable algorithm $\mathcal{A}$ for detection.
While inference for $T$ is intrinsically interesting and important, a side benefit of confidence sets for $T$ is that one can use these to derive post-hoc confidence sets for unknown parameters of $F_{0}$ and $F_{1}$. To clarify, suppose $\mathcal{P}_{i}=\{F_{\theta}\}_{\theta\in\Theta_{i}};i=0,1$. Our method that delivers (3) also delivers confidence sets $\mathcal{C}^{\prime}\subseteq\Theta_{0}\times\Theta_{1}$ such that $\mathbb{P}_{\theta_{0},T,\theta_{1}}((\theta_{0},\theta_{1})\in\mathcal{C}^{\prime}\mid\tau\geq T)\geq 1-\alpha,\forall\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},$ where $\mathbb{P}_{\theta_{0},T,\theta_{1}}$ is the distribution of the sequence $X_{1},\dots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{0}}$, $X_{T},X_{T+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{1}}$. We stress that our confidence sets do not rely on asymptotics and are valid in finite samples.
@@ 1.1 Related work
We are aware of only two directly related works: [ding2003lower] and [wu2007inference, Chapter 3]. They also chose to aim for the guarantee (3) like us. Both [ding2003lower, wu2007inference] provide asymptotically valid inference after the CUSUM procedure [page1954continuous] declares a change for known $F_{0}$ and $F_{1}$ in an exponential family; the former only constructs a lower confidence bound for $T$ while the latter also provides a confidence set for $T$. Their approaches have notable methodological and theoretical limitations: (i) they are tailored specifically to the CUSUM procedure, (ii) they assume that the pre- and post-change distributions are both known and they belong to an exponential family, (iii) they only provide asymptotic coverage guarantees (valid when both $T\to\infty$ and ARL $\to\infty$). These constraints limit practical applicability, and their confidence set exhibits undercoverage in our experiments.
In sharp contrast, our proposed confidence sets for $T$ are very general: We make no assumptions on the detection procedure $\mathcal{A}$, and our framework acts as a versatile wrapper that can be layered on top of any sequential change detection algorithm. It is not restricted to exponential families — it works for any known $F_{0}$ and any nonparametric post-change class $\mathcal{P}_{1}$ on any general space (e.g., $\mathbb{R}^{d}$) and offers non-asymptotic coverage guarantees. It also extends to composite pre-change classes under some assumptions. Thus, our framework is the first nonparametric, theoretically sound, and widely applicable practical tool for sequential change localization tasks. We summarize the comparison in Table 1.
Table 1: Comparison of existing methods and our proposed approach
Pre-, Post- Existing Methods [ding2003lower, wu2007inference] Our Proposed Method
@@
($F_{0},F_{1}$) $\in$ exponential family ($T$, ARL $\to\infty$) (any $\mathcal{A},F_{0},F_{1}$) (Theorem 6)
@@
($F_{0},\mathcal{P}_{1}$) (any $\mathcal{A},F_{0},\mathcal{P}_{1}$) (Theorem 2)
Unknown, Unknown No existing work Assumption 1 Nonasymptotic
($\mathcal{P}_{0},\mathcal{P}_{1}$) on $\mathcal{A},\mathcal{P}_{0}$ (any $\mathcal{P}_{1}$) (Theorem 4)
@@
Literature on sequential change detection. In sequential change detection, the goal is to detect that a change has occurred as soon as possible, but not to provide inferential tools for when the change occurred. This methodology has proved crucial in many applications, such as quality control, cybersecurity, medical diagnostics, and so on. Classical methods such as Shewhart control charts [shewhart1925application], the CUSUM procedure [page1954continuous], and the Shiryaev-Roberts approach [shiryaev1963optimum, roberts1966comparison] assume known pre- and post-change distributions. To address scenarios with unknown post-change distributions, generalized likelihood ratio (GLR) procedures [lorden1971procedures, siegmund1995using] were developed. Comprehensive reviews of these methods are provided by [tartakovsky2014sequential]. More recently, kernel-based techniques [harchaoui2008kernel], e-detectors [shin2022detectors], etc., have emerged to handle nonparametric settings. All these methods focus on the quickest change detection, with average run length (ARL) control. In contrast, another strand of the literature emphasizes controlling the false alarm rate instead of ARL, often at the cost of a longer detection delay [chu1996monitoring, kirch2022sequential, aue2024state]. These approaches typically rely on an initial historical period of data to calibrate the detection thresholds.
@@
Offline vs. online change localization. As far as we are aware, most existing approaches to changepoint localization focus on offline settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times. Classical works, e.g., [hinkley1970inferencce] study the asymptotic behavior of the MLE for the changepoint and construct likelihood-ratio-based confidence sets. [darkhovskh1976a] proposed some nonparametric estimators. [worsley1986confidence] developed likelihood ratio-based confidence sets for exponential family distributions, while [dumbgen1991asymptotic] used bootstrap-based tests for each candidate time to construct nonparametric confidence sets with asymptotic coverage. Recently, [verzelen2023optimal] studies optimal rates for both changepoint detection and localization. [jang2024fast] constructs confidence sets by evaluating a local test statistic. To re-emphasize, all these methods are for the offline setting. In the sequential setting, [Srivastava01011999, Gombay10012003, wu2004bias, Brodsky30042010] develop point estimators for $T$; but very few works construct confidence sets [ding2003lower, wu2007inference], whose drawbacks were discussed earlier.
@@
Inference on $\theta_{0},\theta_{1}$ after stopping. As with the problem of post-detection inference on the changepoint $T$, there is limited literature on valid estimation or inference for the pre- and post-change parameters $\theta_{0}$ and $\theta_{1}$ post-stopping. [WU20063625] constructs asymptotic bias-corrected point estimates and confidence intervals for the post-change mean after a mean change is detected by a CUSUM procedure. [wu2005inference] generalizes these results for the case when there is a possible change in the variance as well. These methods are also discussed in [wu2007inference]. However, these approaches share several limitations with the existing post-detection changepoint inference methods, as mentioned earlier: they rely on asymptotics, assume specific parametric forms, and apply only to the CUSUM procedure.
@@ 1.2 Contributions and paper outline
  • We propose a general framework for constructing a confidence set for the changepoint based on inverting tests obtained using a universal threshold, assuming all observations to be independent. We show that it provides $1-\alpha$ conditional coverage for known pre-change and arbitrary (nonparametric) post-change classes under no additional assumption in Theorem 2. We also extend it for composite pre-change under a suitable assumption in Theorem 4.
  • We propose an alternative method of constructing the tests using adaptive thresholds (which is often tighter than the universal threshold), assuming pre- and post-change distributions to be known, in Algorithm 1. We extend this framework to handle parametric pre- and/or post-change classes in Algorithms 2 and 3. These adaptive methods allow the observations to be dependent. We establish corresponding coverage guarantees in Theorems 6, 7 and 8.
  • As a byproduct of our methods, we derive confidence sets for the unknown pre- and post-change parameters and show their coverage guarantees in Theorem 9.
@@
The rest of the paper is organized as follows. In Section 2, we establish some fundamental results showing the impossibility of unconditional coverage and linking conditional confidence sets to hypothesis tests. Section 3 develops a general framework for constructing a confidence set for $T$ based on inverting tests obtained using a universal threshold. In Section 4, we introduce a simulation-based approach for constructing confidence sets using an adaptive threshold in the parametric setting and study their conditional coverage probabilities. In Section 5, we derive confidence sets for the changepoint, as well as the pre- and post-change parameters. Section 6 presents an extensive simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach  [^2] Codes are available at https://github.com/Aytijhya/Seq-changepoint-localization. This article is concluded in Section 7. All the mathematical proofs, implementation details of some algorithms in the paper, and additional experiments are presented in the Supplementary Material.
@@ 2 Some Fundamental Results
In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$. We begin by clarifying some notational conventions that will be used throughout the paper.
Notation. For $t\in\mathbb{N}$, let $\mathbb{P}_{F_{0},t,F_{1}}$ denote joint distribution of the sequence $\{X_{n}\}_{n}$ when $X_{1},\dots,X_{t-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$ and $X_{t},X_{t+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$.  [^3] To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions. However, our framework in Section 4 can actually handle dependent data. Under dependence, $F_{0}$ and $F_{1}$ denotes the joint distributions of $X_{1},\dots,X_{t-1}$ and $X_{t},X_{t+1},\dots$ respectively, and is in fact ambivalent to whether $X_{t}$ is drawn independently of $X_{1},\dots,X_{t-1}$ or conditional on it, so we do not emphasize the distinction in our notation. As a corner case, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$. In parametric settings, we write $\mathbb{P}_{\theta_{0},t,\theta_{1}}$ and $\mathbb{P}_{\theta_{0},\infty}$ as shorthand for $\mathbb{P}_{F_{\theta_{0}},t,F_{\theta_{1}}}$ and $\mathbb{P}_{F_{\theta_{0}},\infty}$ respectively.
@@
The proposition says that for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$. Thus, for post-detection inference, the usual notion of coverage is not an appropriate one. Instead, we focus on conditional coverage (3), a more reasonable criterion, which was also considered in prior work [ding2003lower, wu2007inference]. The key idea behind our methodology is that for each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (4)
@@
$\mathcal{C}:=\{t\in\mathbb{N}:t\leq\tau,H_{0,t}\text{ is not rejected}\}.$ (5)
@@
Proposition 2.
Let $\tau$ be a stopping time and $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying (7). Then, the test defined by $\phi^{(t)}(X_{1},\cdots,X_{\tau})=\mathds{1}(t\notin\mathcal{C},t\leq\tau)$ satisfies
$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (6)
Conversely, given any test $\phi^{(t)}$ for $H_{0,t}$ using data up to $\tau$ satisfying (6), the inverted confidence set $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\penalty 10000\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (7)